Two cars leave from New York and drive in opposite directions (north and south). The car going north travels at 60 mph. The car going south leaves an hour and a half after the first and goes 80 mph. At what time will the cars be 620 miles apart? This is a homework problem on a sheet of paper so not from a book. I tried it and got 32 hours 30 minutes but I don't think that it's right.
Answer by checkley77(12844)

LET 60*T= THE NORTH CAR.
LET 80(T-1.5)= THE SOUTH CAR.
60T+80(T-1.5)=620
60T+80T-120=620
140T=620+120
140T=740
T=740/140
T=5.2857 HOURS AFTER THE CAR GOING NORTH STARTED THEY WILL BE 620 MILES APART
PROOF:
60*5.2857+80(5.2857-1.5)=620
317.14285+80*3.7857=620
317.14285+302.856=620
620~620